# LLaVA-Mini: Efficient Image and Video Large Multimodal Models with One Vision Token

### v1

@@ 1 Introduction
The computational demands of LMMs are primarily driven by model scale and the number of tokens in the input context. Existing approaches to improving LMM efficiency typically focus on model downsizing (Chu et al., 2023; 2024; Yuan et al., 2024a; Zhou et al., 2024a) or quantization techniques (Yuan et al., 2024b), but often overlook another critical avenue: reducing the number of vision tokens to shorten the input context. Some token reduction methods rely on predefined rules to reduce the number of tokens output by the vision encoder (Bolya et al., 2023; Shang et al., 2024; Li et al., 2024e; Ye et al., 2024c; Hu et al., 2024), which leads to the loss of visual information and inevitably results in performance degradation (Wang et al., 2024; Fan et al., 2024).
@@ 3.2 Preliminary Analyses
Vision Tokens are More Important in Early Layers To find out which layers in LMM the vision tokens play a more important role, we measure the attention weights assigned to different token types (including instruction, vision, and response) at each layer. As shown in Figure 2, the attention assigned to vision tokens varies significantly across layers. Visual tokens receive more attention in the earlier layers, but this attention sharply decreases in the deeper layers, with over 80% of the attention being directed towards instruction tokens. This change in attention suggests that vision tokens play a central role in the early layers, with the instruction tokens seeking relevant visual information from vision tokens through attention mechanisms. In the later layers, the model relies more on instructions that have already fused the visual data to generate responses.
@@
To intuitively illustrate the layer-wise variations in the importance of visual tokens, Figure 4 visualizes the attention distribution across each layer of LLaVA-v1.5. Almost all visual tokens receive broader attention in the early layers, while only some visual tokens are focused in the later layers. These observations suggest that all visual tokens are crucial in the early layers, and reducing their quantity inevitably results in a loss of visual information. This explains why previous methods of direct token reduction will compromise visual understanding capabilities (Shang et al., 2024; Ye et al., 2024c; Hu et al., 2024).
@@ 4.1 Architecture
Vision Token Compression To enhance the efficiency of LMMs, LLaVA-Mini reduces the number of vision tokens fed into the LLM backbone by utilizing a query-based compression module. To learn compression of the vision tokens, LLaVA-Mini introduces $C\times C$ learnable compression queries $\mathbf{Q}^{\textrm{v}}$. These queries interact with all vision tokens $\mathbf{H}^{\textrm{v}}$ through cross-attention (Li et al., 2023a), selectively extracting the important visual information to produce $C\times C$ compressed vision tokens $\mathbf{\hat{H}}^{\textrm{v}}\in\mathbb{R}^{C^{2}\times d_{h}}$. To preserve the spatial information in the image during compression, we introduce a 2D sinusoidal positional encoding $PE(\cdot)$ (He et al., 2021) on the learnable queries and original vision tokens. Formally, the compression can be expressed as:
@@ 5.2 Main Results
Image-based Evaluation We compare LLaVA-Mini with LLaVA-v1.5 across 11 benchmarks to thoroughly assess the performance of LLaVA-Mini with minimal vision tokens. The results are reported in Table 1, where LLaVA-Mini achieves performance comparable to LLaVA-v1.5 while using only 1 vision token instead of 576. Previous efficient LMMs with fewer vision tokens often merged similar tokens directly after the vision encoder (Shang et al., 2024; Ye et al., 2024c), resulting in a significant loss of visual information and negatively impacting visual understanding of LMMs. For instance, LLaMA-VID, VoCo-LLaVA, and MQT-LLaVA, which reduce vision tokens to 1-2 tokens, lead to 5% performance drop on average. In contrast, LLaVA-Mini employs modality pre-fusion to integrate visual information into text tokens before compressing the vision tokens, achieving performance comparable to LLaVA-v1.5 at a token compression rate of 0.17%. Furthermore, LLaVA-Mini-HD shows an average performance improvement of 2.4% over LLaVA-v1.5 due to high-resolution image modeling. Note that LLaVA-Mini-HD has a computational load of 8.13 TFLOPs, which remains lower than LLaVA-v1.5’s 8.55 TFLOPs. More efficiency analyses refer to Sec.5.3. Overall, LLaVA-Mini preserves strong visual understanding capabilities while compressing vision tokens, enhancing the usability of efficient LMMs in visual scenarios.
@@ 6.1 Superiority of Modality Pre-fusion
The proposed modality pre-fusion is central to LLaVA-Mini, as it integrates visual information into text tokens in advance, facilitating extreme compression of vision tokens. To investigate the effects of modality pre-fusion, we conduct an ablation study in Table 7. Without pre-fusion, token compression results in a performance drop of around 5%, even with 144 vision tokens retained, the performance of LMMs falls short of LLaVA-v1.5. This also explains why previous token merging methods often exhibit poor performance (Ye et al., 2024c) or can only achieve a compression rate of over 40% (Shang et al., 2024). Notably, under the same FLOPs, increasing the number of pre-fusion layers yields greater benefits than increasing the number of compression vision tokens. This supports our preliminary analysis, which indicated that vision tokens exhibit varying importance across different layers and vision tokens are more critical in early layers. Investing more computational overhead in earlier stages where vision tokens are more important results in better performance.
@@ D.1 Image-based LMMs
We compare LLaVA-Mini with LLaVA-v1.5 (Liu et al., 2023b) and other advanced LMMs of similar data and model scales, including BLIP-2 (Li et al., 2023a), InstructBLIP (Liu et al., 2024a), IDEFICS (Laurençon et al., 2023), Qwen-VL (Bai et al., 2023), Qwen-VL-Chat (Bai et al., 2023), SPHINX (Lin et al., 2023b), mPLUG-Owl2 (Ye et al., 2024b).
@@
VoCo-LLaMA (Ye et al., 2024c) compresses all vision tokens using language models, significantly improving computational efficiency.
@@ D.2 Video-based LMMs
LLaVA-Mini can also perform high-quality video understanding, so we compare LLaVA-Mini with the current advanced video LMMs, including LLaMA-Adaptor (Zhang et al., 2024), InternVideo (Wang et al., 2022), VideoChat (Li et al., 2024c), Video-LLaMA (Zhang et al., 2023), mPLUG-Owl (Ye et al., 2024a), Video-ChatGPT (Maaz et al., 2024), BT-Adapor (Liu et al., 2023c), LLaMA-VID (Li et al., 2023b), and Video-LLaVA (Lin et al., 2023a).

### v2

@@ 1 Introduction
The computational demands of LMMs are primarily driven by model scale and the number of tokens in the input context. Existing approaches to improving LMM efficiency typically focus on model downsizing (Chu et al., 2023; 2024; Yuan et al., 2024a; Zhou et al., 2024a) or quantization techniques (Yuan et al., 2024b), but often overlook another critical avenue: reducing the number of vision tokens to shorten the input context. Some token reduction methods rely on predefined rules to reduce the number of tokens output by the vision encoder (Bolya et al., 2023; Shang et al., 2024; Li et al., 2024e; Ye et al., 2024d; Hu et al., 2024), which leads to the loss of visual information and inevitably results in performance degradation (Wang et al., 2024; Fan et al., 2024).
@@ 3.2 Preliminary Analyses
Vision Tokens are More Important in Early Layers To find out which layers in LMM the vision tokens play a more important role, we measure the attention weights assigned to different token types (including instruction, vision, and response) at each layer. As shown in Figure 2, Visual tokens receive more attention in the earlier layers, but this attention sharply decreases in the deeper layers, with over 80% of the attention being directed towards instruction tokens. This finding is consistent with the previous conclusion (Chen et al., 2024). This change in attention suggests that vision tokens play a central role in the early layers, with the instruction tokens seeking relevant visual information from vision tokens through attention mechanisms. In the later layers, the model relies more on instructions that have already fused the visual data to generate responses.
@@
To intuitively illustrate the layer-wise variations in the importance of visual tokens, Figure 4 visualizes the attention distribution across each layer of LLaVA-v1.5. Almost all visual tokens receive broader attention in the early layers, while only some visual tokens are focused in the later layers. These observations suggest that all visual tokens are crucial in the early layers, and reducing their quantity inevitably results in a loss of visual information. This explains why previous methods of direct token reduction will compromise visual understanding capabilities (Shang et al., 2024; Ye et al., 2024d; Hu et al., 2024).
@@ 4.1 Architecture
Vision Token Compression To enhance the efficiency of LMMs, LLaVA-Mini reduces the number of vision tokens fed into the LLM backbone by utilizing a query-based compression module. To learn compression of the vision tokens, LLaVA-Mini introduces $C\times C$ learnable compression queries $\mathbf{Q}^{\textrm{v}}$. These queries interact with all vision tokens $\mathbf{H}^{\textrm{v}}$ through cross-attention (Li et al., 2023a; Ye et al., 2024a), selectively extracting the important visual information to produce $C\times C$ compressed vision tokens $\mathbf{\hat{H}}^{\textrm{v}}\in\mathbb{R}^{C^{2}\times d_{h}}$. To preserve the spatial information in the image during compression, we introduce a 2D sinusoidal positional encoding $PE(\cdot)$ (He et al., 2021) on the learnable queries and original vision tokens. Formally, the compression can be expressed as:
@@ 5.2 Main Results
Image-based Evaluation We compare LLaVA-Mini with LLaVA-v1.5 across 11 benchmarks to thoroughly assess the performance of LLaVA-Mini with minimal vision tokens. The results are reported in Table 1, where LLaVA-Mini achieves performance comparable to LLaVA-v1.5 while using only 1 vision token instead of 576. Previous efficient LMMs with fewer vision tokens often merged similar tokens directly after the vision encoder (Shang et al., 2024; Ye et al., 2024d), resulting in a significant loss of visual information and negatively impacting visual understanding of LMMs. For instance, LLaMA-VID, VoCo-LLaVA, and MQT-LLaVA, which reduce vision tokens to 1-2 tokens, lead to 5% performance drop on average. In contrast, LLaVA-Mini employs modality pre-fusion to integrate visual information into text tokens before compressing the vision tokens, achieving performance comparable to LLaVA-v1.5 at a token compression rate of 0.17%. Furthermore, LLaVA-Mini-HD shows an average performance improvement of 2.4% over LLaVA-v1.5 due to high-resolution image modeling. Note that LLaVA-Mini-HD has a computational load of 8.13 TFLOPs, which remains lower than LLaVA-v1.5’s 8.55 TFLOPs. More efficiency analyses refer to Sec.5.3. Overall, LLaVA-Mini preserves strong visual understanding capabilities while compressing vision tokens, enhancing the usability of efficient LMMs in visual scenarios.
@@ 6.1 Superiority of Modality Pre-fusion
The proposed modality pre-fusion is central to LLaVA-Mini, as it integrates visual information into text tokens in advance, facilitating extreme compression of vision tokens. To investigate the effects of modality pre-fusion, we conduct an ablation study in Table 7. Without pre-fusion, token compression results in a performance drop of around 5%, even with 144 vision tokens retained, the performance of LMMs falls short of LLaVA-v1.5. This also explains why previous token merging methods often exhibit poor performance (Ye et al., 2024d) or can only achieve a compression rate of over 40% (Shang et al., 2024). Notably, under the same FLOPs, increasing the number of pre-fusion layers yields greater benefits than increasing the number of compression vision tokens. This supports our preliminary analysis, which indicated that vision tokens exhibit varying importance across different layers and vision tokens are more critical in early layers. Investing more computational overhead in earlier stages where vision tokens are more important results in better performance.
@@ 7 Conclusion
Acknowledgments
We thank all the anonymous reviewers for their insightful and valuable comments. This work was supported by National Natural Science Foundation of China (Grant No.62376260).
@@ D.1 Image-based LMMs
We compare LLaVA-Mini with LLaVA-v1.5 (Liu et al., 2023b) and other advanced LMMs of similar data and model scales, including BLIP-2 (Li et al., 2023a), InstructBLIP (Liu et al., 2024a), IDEFICS (Laurençon et al., 2023), Qwen-VL (Bai et al., 2023), Qwen-VL-Chat (Bai et al., 2023), SPHINX (Lin et al., 2023b), mPLUG-Owl2 (Ye et al., 2024c).
@@
VoCo-LLaMA (Ye et al., 2024d) compresses all vision tokens using language models, significantly improving computational efficiency.
@@ D.2 Video-based LMMs
LLaVA-Mini can also perform high-quality video understanding, so we compare LLaVA-Mini with the current advanced video LMMs, including LLaMA-Adaptor (Zhang et al., 2024), InternVideo (Wang et al., 2022), VideoChat (Li et al., 2024c), Video-LLaMA (Zhang et al., 2023), mPLUG-Owl (Ye et al., 2024b), Video-ChatGPT (Maaz et al., 2024), BT-Adapor (Liu et al., 2023c), LLaMA-VID (Li et al., 2023b), and Video-LLaVA (Lin et al., 2023a).